OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks.Train A is traveling at 40 miles per hour and train B is traveling at 60 miles per hour. Train A passes a station at 3:25am. If train B passes the same station at 3:37am, at what time will train B catch up to train A?

OpenStudy (saifoo.khan):

The problem starts at the station. B leaves the station 12 minutes, or (1/5) hr., after A. When B catches A they are the same distance from the station. ------------------ DATA 1; rate = 40 mph ; time = x hrs ; distance = 40x miles DATA 2: rate = 60 mph ; time = (x-(1/5)) ; distance = 60(x-(1/5)) = 60x-12 miles EQUATION: 40x = 60x-12 20x = 12 x = 3/5 hr = 36 minutes 3:25 + 36 = 4:01 AM (time A catches up to B)